What is the molality of a solution composed of 40.0 g sodium nitrate (NaNO3) and 500.0 g of water? 


1
Expert's answer
2022-04-01T02:58:04-0400

Solution:

Molality = Moles of solute / Kilograms of solvent


solute = NaNO3

solvent = H2O (water)


The molar mass of NaNO3 is 84.995 g/mol

Therefore,

Moles of NaNO3 = (40.0 g NaNO3) × (1 mol NaNO3 / 84.995 g NaNO3) = 0.4706 mol NaNO3

Moles of NaNO3 = 0.4706 mol


Kilograms of water = (500.0 g H2O) × (1 kg / 1000 g) = 0.5 kg H2O


Thus:

Molality of NaNO3 solution = Moles of NaNO3 / Kilograms of water

Molality of NaNO3 solution = (0.4706 mol) / (0.5 kg) = 0.94 mol/kg = 0.94 m

Molality of NaNO3 solution = 0.94 m


Answer: The molality of NaNO3 solution is 0.94 m